The maximum value of cos^2 (pie/4-x)+(sinx- cosx)^2 is

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
maximum value of cos^2 (pie/4-x)+(sinx- cosx)^2 isCos2A= 2cos^2A - 1So Cos^2 (pie/4-x) = cos ( pi/2-2x( = sin2xHere cos^2 (pie/4-x)+(sinx- cosx)^2 =Sin2X+ sin^x+ cos^2x- 2sincosx=1+sin2x- Sin2X= 1